# Evaluating the impact of racial disproportionality in Stop & Search on expressive crimes in London

### v1

@@ Abstract
Racial disproportionality in Stop & Search practices elicits substantial concerns about its societal and behavioral impacts. This paper aims to investigate the effect of disproportionality, particularly on the black community, on expressive crimes in London using data from January 2019 to December 2023. We focus on a semi-parametric partially linear structural regression method and introduce a Bayesian empirical likelihood procedure combined with double machine learning techniques to control for high-dimensional confounding and to accommodate the strong prior assumption. In addition, we show that the proposed procedure generates a valid posterior in terms of coverage. Applying this approach to the Stop & Search dataset, we find that racial disproportionality aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on expressive crimes.
Keywords: Stop & Search; Racial disproportionality; Partially linear regression; Bayesian estimation; Double machining methods
@@ 1.1 Backgroud
Racial disproportionality in Stop & Search practices, particularly in cities such as London, raises important questions about its broader social and behavioral effects. Expressive crimes, such as vandalism, public disorder, or gang-related violence, are often argued to be influenced by social alienation, perceived injustice, or strained community-police relations, which can result from disproportionate policing practices.
@@
Bowling and Phillips (2007) suggested that racial discrimination within policing is exacerbated by the Stop & Search powers in the UK targeting Black individuals disproportionately. The UK’s State of Policing report in 2022 highlighted that Stop & Search is an important tool for preventing and detecting crime, with significant public support when used fairly and proportionately, particularly in targeting weapons and drugs (HM Inspectorate of Constabulary and Fire & Rescue Services, 2022). Tiratelli et al. (2018) assessed the effectiveness of stop-and-search in reducing crime in London, with findings indicating some correlation between stop-and-search and drug offenses, but relatively limited impact on broader crime reduction.
@@
In recent years, the issue of disproportionality in Stop & Search practices has remained a significant concern in London, where the disparities between ethnic groups have persisted despite various reform efforts. The March 2024 Disproportionality Board Data Pack (MOPAC, 2023), part of the Mayor’s Action Plan for Transparency, Accountability, and Trust in Policing, highlighted that Black individuals in London are still 3.3 times more likely to be stopped and searched compared to White individuals, and this figure rises to 6.2 times for searches related to weapons.
@@
Scholars have identified that Stop & Search are highly concentrated in low-income neighborhoods or areas with higher populations of minority ethnic groups. According to Millner (2020), the increased use of new surveillance technologies, such as predictive policing software, has transformed policing in cities such as London, intensifying the monitoring of particular demographic groups. Suss and Oliveira (2023) and Meng (2017) demonstrated that the spatial patterns of Stop & Search in London are closely linked to areas characterized by higher levels of economic inequality and minority populations, revealing racial bias embedded in police practices. Oberwittler and Roché (2022) argued that in France, Germany and other European cities, police actions against adolescents in certain neighborhoods are often shaped by institutional biases tied to economic deprivation and race. Given these findings, it is crucial to use a data-driven approach to quantify the effect of racial disproportionality in Stop & Search. For example, a robust statistical approach is needed to account for the role of various confounding factors, such as differences in demographics across communities, socioeconomic conditions and policing priorities, from potential biases in police decision-making.
1.2 Statistical challenges
@@
Based on the data from January 2019 to December 2023 in London aggregated within the borough level, the objective of this study is to estimate the effect of disproportionality in Stop & Search on expressive crimes. Estimation of the effect of disproportionality in Black people is complicated by the potential for confounding between the proportion of black population and the outcome. Confounding exists whenever exposure (or treatment) assignment is dependent on predictors that also influence the outcome, and appropriate adjustment is required to estimate the corresponding effect. Variables, such as the number of schools and proportion of the green space, can obscure the true relationship between these variables. For example, higher Stop & Search rates for Black people in areas might be attributed to an increased proportion of the black population in this borough, but it could also be influenced by other factors like the proportion of the educated household and the number of schools in the area. Determining this effect in this complex scenario requires careful consideration of variables and the implementation of rigorous research in data analysis. To estimate the effect of a policy, a common approach is to use partially linear regression as in Robinson (1988), which involves specifying separate regression models for both the treatment and outcome variables. The model for the treatment variable is often termed as ‘propensity score’ model. Propensity score adjustments (Rosenbaum and Rubin, 1983) have been extensively used to reduce confounding bias in estimating causal effects. The propensity score is defined as the conditional probability of the treatment assignment given confounding covariates. The propensity score can be used to break the dependence between confounders and exposure, to create balance in the distribution of confounders across exposure groups, and to facilitate correct inference. However, a key statistical challenge in applying partially linear regression for Stop & Search practices is the presence of high-dimensional confounders, where the parameter space for nuisance parameters grows with the sample size. In this case, the traditional semi-parameter theory might not offer valid inference for the parameter of interest. To overcome this challenge, Chernozhukov et al. (2018) proposed the use of machine learning methods to estimate the nuisance parameter and provided a simple and root-$n$ consistent procedure to estimate the parameter of interest via the Neyman moment equation and sample splitting.
@@
Another challenge arises, when analyzing Stop & Search data, due to the strong prior beliefs and assumptions often held by policymakers and law enforcement agencies. These priors may reflect institutional perspectives on crime prevention strategies or operational practices, and we need to account for these prior beliefs in order to ensure robust and objective analysis, fostering evidence-based policy decisions. Therefore, a prior-to-posterior Bayesian analysis is preferable. The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the effect from a practical perspective. There is growing interest in the application of Bayesian methodology to two-stage propensity score regression analysis; however, most of existing Bayesian approaches require a full parametric specification for both structural equations in both regression models (see, for example McCandless et al., 2010; Kaplan and Chen, 2012). There are several works already attempting to solve this problem via a semi-parameter perspective (for example, Graham et al., 2016; Liu et al., 2020); these methods typically exploit the Bayesian bootstrap (Rubin, 1981; Chamberlain and Imbens, 2003) to perform inference. Recently, Luo et al. (2023) proposed to draw inference for the posterior from a Bayesian predictive distribution via a Dirichlet process model, extending the Bayesian bootstrap, and opening up the possibility to explore the new direction on doubly robust causal inference based on a non-parametric specification. In their work, the posterior samples are generated by resampling weights from the Pólya urn. It links with recent advancements in Bayesian empirical titled likelihood (Chib et al., 2018; Yiu et al., 2020; Luo et al., 2023), where the weights are replaced by the empirical probabilities. Antonelli et al. (2022) proposed to use the Gaussian process (GP) regression, and then the MCMC estimate is plugged into estimating equation. However, in present of high-dimensional confouders, those approaches can be slow when implementing Markov chain Monte Carlo (MCMC). In this paper, we aim to marry between the Bayesian method and the semi-parametric double machine learning method and provide a computationally efficient procedure to draw inference on the parameter of interest. We argue that our method bridges this gap by demonstrating how the approximate frequentist distribution theory can find an equally effective interpretation within a Bayesian framework in the high-dimensional setting. In addition, we show that the proposed procedure generates a valid posterior according to Monahan and Boos (1992), indicating a valid putative ‘posterior’ density computed by a non-standard method should still make probability statements consistent with Bayes’ rule.
1.3 Plan of paper
@@
The remainder of this paper is organized as follows. In Section 2, we review the estimation procedure of partially linear regression, and introduce the notion of approximating Bayesian formulations and how to perform inference via the Neyman moment equation. We verify the validity of the proposed posterior inference in the spirit of Monahan and Boos (1992) in Section 2.3. Section 3 shows some simulation examples which compare the proposed method with some other frequentist and Bayesian approaches, following with Stop & Search data analysis in Section 4. Finally, Section 5 presents some concluding remarks and future research directions.
2 Statistical methodology
@@
To estimate the effect of racial dispropotionality, we first consider the partially linear regression in Robinson (1988):
$Y=\mu(X)+\beta D+U,\;\;\;\mathbb{E}(U|X,D)=0,\;\;\;\mathbb{E}(D|X)=\pi(X)\$ (1)
where $Y$ is the outcome variable, $D$ is the treatment variable of interest (i.e., the proportion of Black people), and $X=\left(X_{1},\ldots,X_{p}\right)$ is a list of counfounders. Given independent and identically distributed data, $\left\{Z_{i}=(Y_{i},D_{i},X_{i}),i=1,\ldots,n\right\}$, the parameter of interest is $\beta$. If $D$ is exogenous conditional on $X$, then $\beta$ has the interpretation of the treatment effect parameter. Assuming there are no unmeasured confounders Under frequentist inference, the estimand of $\beta$ is consistent if $\mathbb{E}\left[Y\left|X,D\right.\right]$ is correctly specified. In addition, one can be constructed via specifying both outcome mean and propensity score models and use them in a combined estimation to yield a debiased estimator. To implement such inference, we need to specify the following two models:
@@
  • Propensity score (PS) model: $\mathbb{E}(D|X)$, and it represents the conditional distribution of treatment given the confounders;
  • Outcome regression (OR) model: $\mathbb{E}\left(Y\left|X,D\right.\right)$, which represents the conditional distribution of the outcome given the treatment variable and confounder under the observational study.
@@
$Y-\mathbb{E}[Y|\pi(X)]=\beta[D-\pi(X)]+V,\;\;\;V=U+\mu(X)-\mathbb{E}[\mu(X)|\pi(X)]=0,\;\;\;V\perp\!\!\!\perp D|\pi(X)\$ (2)
and therefore $\mathbb{E}(V|\pi(X))=0$. The treatment effect parameter, $\beta$, can be estimated via the least square estimation procedure. This is equivalent to the Neyman-orthogonal moment equation noted in Chernozhukov et al. (2018), and $\beta$, is the solution $\mathbb{E}[\psi(Z;\beta)]=0$, where
@@
In this way, we can reduce the high-dimensional problem to a single parameter problem if we assume $D$ is a scalar. This approach requires that we have access to both $\mu(X)$ and $\pi(X)$. The challenge arises in the high-dimensional setting, i.e., $p>n$; however, both of these functions can be estimated via machine learning (ML) methods. This double ML specification gives us flexibility to specify the structures of $\mu(X)$ and $\pi(X)$ while focusing on the treatment effect. However, it would be difficult to perform conventional Bayesian analysis with an strong prior input from the policy makers as there is in no distributional form for the conditional outcome mean and PS models. Conventional Bayesian inference focuses on updating prior belief in light of the data, and the data are summarized in the form of the likelihood. The relationship between prior beliefs and observable random quantities, $\mathbf{z}=\left(z_{1},\ldots,z_{n}\right)$, is formulated via the de Finetti representation, i.e.,
@@
2.1 Bayesian non-parametric procedure for the Neyman moment equation
[…]

### v2

# A scalable Bayesian double machine learning framework for high dimensional causal estimation, with application to racial disproportionality assessment

@@ Abstract
Racial disproportionality in Stop and Search practices elicits substantial concerns about its societal and behavioral impacts. This paper aims to investigate the effect of disproportionality, particularly on the black community, on expressive crimes in London using data from January 2019 to December 2023. We focus on a semi-parametric partially linear structural regression method and introduce a scalable Bayesian empirical likelihood procedure combined with double machine learning techniques to control for high-dimensional confounding and to accommodate strong prior assumptions. In addition, we show that the proposed procedure yields a valid posterior in terms of coverage. Applying this approach to the Stop and Search dataset, we find that racial disproportionality aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on expressive crimes.
Keywords: Partially linear regression; Bayesian estimation; Double machining methods; Stop and search; Racial disproportionality
@@ 1.1 Background
Racial disproportionality in stop and search practices, particularly in cities such as London, raises important questions about its broader social and behavioral effects. Expressive crimes, such as vandalism, public disorder, or gang-related violence, are often argued to be influenced by social alienation, perceived injustice, or strained community-police relations, which can result from disproportionate policing practices.
@@
Bowling and Phillips (2007) suggested that racial discrimination within policing is exacerbated by the stop and search police powers in the UK targeting Black individuals disproportionately. The UK’s State of Policing report in 2022 highlighted that stop and search is an important tool for preventing and detecting crime, with significant public support when used fairly and proportionately, particularly in targeting weapons and drugs (HM Inspectorate of Constabulary and Fire & Rescue Services, 2022). Tiratelli et al. (2018) assessed the effectiveness of stop and search in reducing crime in London, with findings indicating some correlation between stop and search and drug offenses, but relatively limited impact on broader crime reduction.
@@
In recent years, the issue of disproportionality in stop and search practices has remained a significant concern in London, where the disparities between ethnic groups have persisted despite various reform efforts. The March 2024 Disproportionality Board Data Pack (MOPAC, 2023), part of the Mayor’s Action Plan for Transparency, Accountability, and Trust in Policing, highlighted that Black individuals in London are still 3.3 times more likely to be stopped and searched compared to White individuals, and this figure rises to 6.2 times for searches related to weapons.
@@
Scholars have identified that stop and search events are highly concentrated in low-income neighborhoods or areas with higher populations of minority ethnic groups. According to Millner (2020), the increased use of new surveillance technologies, such as predictive policing software, has transformed policing in cities such as London, intensifying the monitoring of particular demographic groups. Suss and Oliveira (2023) and Meng (2017) demonstrated that the spatial patterns of stop and search in London are closely linked to areas characterized by higher levels of economic inequality and minority populations, revealing racial bias embedded in police practices. Oberwittler and Roché (2022) argued that in France, Germany and other European cities, police actions against adolescents in certain neighborhoods are often shaped by institutional biases tied to economic deprivation and race. Given these findings, it is crucial to use a data-driven approach to quantify the effect of racial disproportionality in Stop & Search. For example, a robust statistical approach is needed to account for the role of various confounding factors, such as differences in demographics across communities, socioeconomic conditions and policing priorities, from potential biases in police decision-making.
1.2 Statistical challenges and related work
Disparities in institutional decision-making have significant implications for both individual well-being and broader societal outcomes. In healthcare, unequal access to treatments or disproportionality in clinical decision-making across population groups can lead to persistent health inequities. Similarly, differential treatment by law enforcement in policing may raises important questions about its broader social and behavioral effects. Quantifying racial bias in policing has become a central question in policy-oriented social science and applied statistics. Despite wide availability of administrative policing data, such as stop records, arrest databases, and officer-reported activity logs, using these data to make credible causal inferences about discrimination remains methodologically challenging. Knox et al. (2020) demonstrate that administrative policing datasets inherently condition on the outcome of police discretion, and show that estimands targeting racial bias that condition on being stopped can potentially have the opposite signs as the estimands of interest. A policing agency could appear unbiased or even favorable to minority groups among stopped individuals while simultaneously discriminating in its choice of whom to stop. Based on this, Zhao et al. (2022) clarify the definition of discrimination in causal inference terms and emphasize the need for estimands that avoid conditioning on post-treatment variables. Their work stresses that identifying racial disparities requires not only conditioning strategies but careful articulation of what causal question is being answered.
From a causal inference perspective, estimation of racial disproportionality is complicated by confounding of the exposure-outcome association. Confounding exists whenever exposure (or treatment) assignment is dependent on predictors that also influence the outcome, and appropriate adjustment is required to estimate the corresponding effect. Variables, such as the number of schools and proportion of green space, can obscure the true relationship between these variables. For example, higher stop and search rates for Black people in areas might be attributed to an increased proportion of the black population in this borough, but it could also be influenced by other factors like the proportion of the educated household and the number of schools in the area. Determining this effect in this complex scenario requires careful consideration of variables and the implementation of rigorous research in data analysis. Recent methodological advances seek to formalize these adjustments. Gaebler et al. (2022) argue that many policing studies lack explicit causal models, leading to implicit identifying assumptions. Jung et al. (2018) propose a risk-adjusted regression framework to mitigate included- and omitted-variable bias and to focus disparity estimates on differences not explained by legitimate police-relevant risk factors. Recently, Huang et al. (2024) introduce a mobility-adjusted causal estimand, demonstrating that individuals’ movement patterns fundamentally shape exposure to policing. Another potential threat to identification in this context is selection bias, since stop and search events often arise because of a form of racial bias. The data we get to observe arise from interactions the police have with the civilians they choose to stop, and are certainly not a random sample of all police-civilian interactions.
@@ 1.2 Statistical challenges and related work
To estimate causal effects in such complex settings, applied researchers often adopt flexible semiparametric models. A commonly used tool is the partially linear regression framework (Robinson, 1988), which specifies separate regression functions for the treatment and outcome. The model for the treatment variable is often termed as ‘propensity score’ model. Propensity score adjustments (Rosenbaum and Rubin, 1983) have been extensively used to reduce confounding bias in estimating causal effects. The propensity score is defined as the conditional probability of receiving treatment given confounding covariates. The propensity score can be used to break the dependence between confounders and exposure, to create balance in the distribution of confounders across exposure groups, and to allow valid causal inference. However, a key statistical challenge in applying partially linear regression for stop and search practices is the presence of high-dimensional confounders, where the parameter space for nuisance parameters grows with the sample size. In this case, traditional semi-parameter theory might not offer valid inference for the parameter of interest. To overcome these challenges, Chernozhukov et al. (2018) proposed the use of machine learning methods to estimate the nuisance parameters and provided a simple and root-$n$ consistent procedure to estimate the parameter of interest via the Neyman moment equation and sample splitting.
@@
Another challenge arises, when analyzing stop and search data, due to the strong prior beliefs and assumptions often held by policymakers and law enforcement agencies. These priors may reflect institutional perspectives on crime prevention strategies or operational practices, but might also be potentially biased or based on anecdotal evidence. A prior-to-posterior Bayesian analysis allows to incorporate beliefs that are based on prior data-driven evidence, fostering evidence-based policy decisions. Further, unlike approaches that rely on large-sample approximations, which may be unreliable in finite samples or when complex, high-dimensional nuisance models are required, Bayesian inference can deliver more stable and valid uncertainty quantification even in small sample sizes that often arise in areal-level analyses. In this work, we aim to incorporate Bayesian inference within a double machine learning framework to obtain coherent probabilistic statements about the effects of disproportionality, enhance the stability of estimation, and allow the incorporation of domain knowledge. This offers a coherent updating framework for regularization and prior information, which are valuable in high-dimensional or weak-signal settings common in health and social policy applications. There is growing interest in the application of Bayesian methodology to two-stage propensity score regression analysis; however, most of existing Bayesian approaches require a full parametric specification for both structural equations in both regression models (see, for example McCandless et al., 2010; Kaplan and Chen, 2012). There are several works already attempting to solve this problem via a semi-parameter perspective (for example, Graham et al., 2016; Liu et al., 2020); these methods typically exploit the Bayesian bootstrap (Rubin, 1981; Chamberlain and Imbens, 2003) to perform inference. Recently, Luo et al. (2023) proposed to draw inference for the posterior from a Bayesian predictive distribution via a Dirichlet process model, extending the Bayesian bootstrap, and opening up the possibility of performing doubly robust causal inference based on a non-parametric specification. In their work, the posterior samples are generated by resampling weights from the Pólya urn. It links with recent advancements in Bayesian empirical titled likelihood (Chib et al., 2018; Yiu et al., 2020; Luo et al., 2023), where the weights are replaced by the empirical probabilities. Antonelli et al. (2022) proposed to use the Gaussian process (GP) regression, and then the MCMC estimate is plugged into estimating equation. However, despite these advances, computational efficiency and theoretical guarantees remain open challenges, where covariates may be high dimensional and the nuisance components complex. In this paper, we aim to marry the Bayesian method and the semi-parametric double machine learning framework and provide a computationally efficient procedure to draw inference on the parameter of interest. We argue that our method bridges this gap by demonstrating how the approximate frequentist distribution theory can find an equally effective interpretation within a Bayesian framework in the high-dimensional setting. In addition, we show that the proposed procedure generates a valid posterior according to Monahan and Boos (1992), indicating a valid putative ‘posterior’ density computed by a non-standard method should still make probability statements consistent with Bayes’ rule.
@@
The remainder of this paper is organized as follows. Section 2 describes the motivating dataset used in this paper. In Section 3, we articulate the causal question for the study and define the associated causal estimand. In addition, we review the estimation procedure of partially linear regression, and introduce the notion of approximating Bayesian formulations and how to perform inference via the Neyman moment equation. We verify the validity of the proposed posterior inference in the spirit of Monahan and Boos (1992) in Section 3.5. Section 4 shows some simulation examples which compare the proposed method with some other frequentist and Bayesian approaches, following with Stop and Search data analysis in Section 5. Finally, Section 6 presents some concluding remarks and future research directions.
2 Motivating example: Racial disproportionality assessment in Stop and Search
The Stop and Search dataset consists of individual stop and search monthly records in London from January 2019 to December 2023, including date and time, street-level location, ethnicity, gender and age of the person stopped, legislation, object searched and outcome. Further description of this dataset can be found in the Appendix.
Based on 2.6 million stop and search incidents from January 2019 to December 2023 in London aggregated within the borough level, the objective of this study is to examine how the level of disproportionality in stop and search practices for expressive crime targeting Black individuals varies across London boroughs. That is, the focus is on borough-level inequality in policing outcomes rather than on the overall presence or absence of racial bias across the city. Even if disproportionality is high across London, the analysis seeks to identify whether certain boroughs exhibit relatively greater or lesser disproportionality than the average level in London.
Therefore, we define the outcome variable as the disproportionality index (DI) for expressive crime targeting Black people in each Borough (total 33 Boroughs) and it can be calculated as:
[…]
This index contrasts a Borough-specific stop and search rate involving Black individuals and the average rate in London, where $\text{DI}=1$ indicates no disproportionality, whereas $\text{DI}<1$ indicates a stop and search rate below expectation while $\text{DI}>1$ indicates a stop and search rate above expectation.
Figure 1: DI for Black People in Stop & Search for Expressive Reasons
Figure 1 illustrates the DI across boroughs in London. The South-West London boroughs of Richmond upon Thames and Kingston upon Thames and the City of London have the largest DI values. These boroughs have considerably larger index values, revealing stop and search rates above expectation among the Black population. In contrast, east end London boroughs, especially the riverside boroughs — Lewisham, Greenwich, Bexley, Newham, and Barking and Dagenham — all have low DI scores. Boroughs nearer the perimeter of the city, like Enfield and Croydon, also have lower DI scores. This trend indicates a gradient in the distribution, with more central and wealthy areas tending to have larger index values and outer or more residential boroughs having comparatively lower scores.
To evaluate the effect of disproportionality, we use the percentage of Black people in each borough as the treatment variable. Specifically, areas with a higher percentage of Black residents might experience more police attention, which could inflate the DI. If policing practices were biased or influenced by racial profiling, the Stop and Search rates for Black individuals might rise disproportionately, driving up the DI. Moreover, in boroughs with larger Black populations, more recorded crimes against Black individuals could naturally occur, but if the rate of police actions exceeds the expected level in London based on population share, it would signal disproportionality. For example, if Black residents make up 30% of a borough but account for 60% of Stop and Search incidents, the DI would be elevated, demonstrating a disproportionate impact.
The post-treatment variable considered in Zhao et al. (2022), police detainment, can be viewed as analogous to the outcome in this paper, since the disproportionality index is constructed from the frequency of “detainments” or stop interactions involving individuals from minority backgrounds. The key distinction is that our analysis relies on areal-level data, whereas Knox et al. (2020); Zhao et al. (2022) work with individual-level observations. Moreover, in our setting, outcomes for minority groups are not directly contrasted with those for non-minority groups. Instead, the potential outcomes for minority individuals correspond to different “doses” of exposure across districts. From a policy perspective, this raises the question of whether such patterns can be interpreted as evidence of potential discrimination.
3 Methodology
3.1 Notation and causal assumptions
Let $\left\{Z_{i}=(Y_{i},D_{i},X_{i}),i=1,\ldots,n\right\}$ denote independent and identically distributed data measured on $n$ study units (e.g., different boroughs), where $Y_{i}$ denotes an observed outcome (e.g., the disproportionality index DI), $D_{i}$ represents an exposure or treatment, potentially continuous (e.g., percentage of Black people in an area), and $X_{i}$ is a vector of potential confounders measured on unit $i$. The primary focus of our inference is on average potential outcomes (APOs) for a population under a hypothetical exposure. We let $Y_{i}(d)$ be the (potential) DI that would be observed in borough $i$ if the proportion of the Black population were set to $d$. The target estimand is the average causal effect of the Black population proportion on stop and search disproportionality, $\beta=\mathbb{E}\left[\frac{\partial Y(d)}{\partial d}\right]$, capturing the average difference in disproportionality. This estimand captures the causal effect of borough-level racial composition on the degree of disproportionality in stop and search incidents involving Black individuals for expressive crimes. Given these confounders, denoted as $X_{i}$, in borough $i$, our framework relies on the usual core causal assumptions: consistency, Stable Unit Treatment Value (SUTVA), no unmeasured confounding (NUC), and positivity (Antonelli et al., 2022; Li et al., 2023). Under these assumptions, the distribution of potential outcomes is identified from observed data via the g-formula: $E\left[Y(d)\right]=\mathbb{E}\left\{\mathbb{E}\left[Y|D=d,X\right]\right\}$. In the next section, we will specifically focus on partially linear regression to estimate this causal effect using a Bayesian double machine learning approach.
3.2 Partially linear regression
@@
To estimate the effect of policy, i.e. racial dispropotionality, we consider the partially linear regression described in Robinson (1988):
$Y=\mu(X)+\beta D+U,\;\;\;\mathbb{E}(U|X,D)=0,\;\;\;\mathbb{E}(D|X)=\pi(X)\\$ (1)
where $Y$ is the outcome variable, $D$ is the treatment variable, $X=\left(X_{1},\ldots,X_{p}\right)$ is a list of counfounders and $U$ represents the unobserved error term.
[…]
@@
  • Propensity score (PS) model: $\mathbb{E}(D|X)$, which represents the conditional distribution of treatment given the confounders;
[…]
@@
$Y-\mathbb{E}[Y|\pi(X)]=\beta[D-\pi(X)]+V,\;\;\;V=U+\mu(X)-\mathbb{E}[\mu(X)|\pi(X)]=0,\;\;\;V\perp\!\!\!\perp D|\pi(X)\\$ (2)
and therefore $\mathbb{E}(V|\pi(X))=0$. This implies that $cov(D-\pi(X),V)=0$, and the treatment effect parameter, $\beta$, can be estimated via the least square estimation procedure from regressing $Y-\mathbb{E}[Y|\pi(X)]$ on $D-\pi(X)$. However, Hahn (1998) demonstrated that the estimator is not semiparametrically efficient when both $\mathbb{E}[Y|\pi(X)]$ and $\pi(X)$ are estimated non-parametrically. To address this, Chernozhukov et al. (2018) proposed estimating $\beta$ using the Neyman-orthogonal moment equation, which achieves the semi-parametric efficiency bound as it protects the estimator against first-order bias from nuisance function estimation errors. That is, $\beta$ is the solution $\mathbb{E}[\psi(Z;\beta)]=0$, where
@@
We can estimate nuisance functions, $\pi(X)$ and $\mu(X)$, non-parametrically or using machine learning (ML) methods regardless of the dimension of $X$. In this way, we can reduce the high-dimensional problem to a single parameter problem if we assume $D$ is a scalar. This approach requires that we have access to both $\mu(X)$ and $\pi(X)$. The challenge arises in the high-dimensional setting, i.e., $p>n$; however, both of these functions can be estimated via ML methods. This double ML specification gives us flexibility to specify the structures of $\mu(X)$ and $\pi(X)$ while focusing on the treatment effect. However, it would be difficult to perform conventional Bayesian analysis with an strong prior input from the policy makers as there is no distribution assumption in this semi-parametric procedure. Conventional Bayesian inference focuses on updating prior belief in light of the data, and the data are summarized in the form of the likelihood. The relationship between prior beliefs and observable random quantities, $\mathbf{z}=\left(z_{1},\ldots,z_{n}\right)$, is formulated via the de Finetti representation, i.e.,
@@
3.3 Bayesian non-parametric procedure for the double machine learning method
Suppose we assume that there is a set of Neyman-orthogonal equations $\psi(Z;\beta)$ such that $\mathbb{E}[\psi(Z;\beta)]=0$, for all $\beta$. The objective is to find a non-parametric approximate likelihood, $p(z)$, to the true data generating model $f(z|\beta)$. Therefore, we can define some discrepancy $\delta(f|p)$, subject to $\int\psi(z;\beta)p(z)dz=0$ and $\int p(z)dz=1$. This becomes an optimization problem:
$\min_{p}\delta(f|p)\text{ subject to }\int\psi(z;\beta)p(z)dz=0\;\;\int p(z)dz=1,\;\;\forall\beta\in\mathbb{R}.$ (3)